CAVITY RINGDOWN SPECTROSCOPY OF CIS-CIS HOONO AND THE HOONO/HONO2 BRANCHING RATIO IN THE REACTION OH + NO2 + M

\underline{ANDREW K. MOLLNER}, BRIAN D. BEAN, SERGEY NIZKORODOV, GAUTHAM NAIR, MIT-CHIO OKUMURA, California Institute of Technology, Pasadena, CA 91125, USA; STANLEY P. SANDER, NASA Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109, USA; KIRK A. PETERSON, Department of Chemistry, Washington state University, Pullman, WA 99164, USA; JOSEPH S. FRANCISCO, Department of Chemistry, Purdue University, West Lafayette, IN 47907, USA.

The termolecular association reaction OH + NO2 + M was studied in a low pressure discharge flow reactor, and both HONO2 and HOONO products were detected by Infra-Red Cavity Ringdown Spectroscopy (IR-CRDS). The absorption spectrum of the fundamental ν1 band of the cis-cis isomer of HOONO (pernitrous or peroxynitrous acid) was observed at 3306 cm^{−1}, in good agreement with matrix isolation studies and ab initio predictions. The rotational contour of this band was partially resolved at 1cm^{−1} resolution and matched the profile predicted by ab initio calculations. The integrated absorbances of the ν1 bands of the cis-cis HOONO and HONO2 products were measured as a function of temperature and pressure.